# On a novel UCP result and its application to inverse conductive scattering

### v2

@@ 3. Unique recovery results for the conductive polygonal-nest and polygonal-cell medium scatterers
$\mathcal{P}_{4}$ $=k^{2}(q_{1}^{(1)}-q_{2}^{(1)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{1}\tilde{\delta}u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}+k^{2}(q_{1}^{(2)}-q_{2}^{(2)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{2}\tilde{\delta}u_{2}^{-}u_{0}\mathrm{d}\mathbf{x},$
@@
$\mathcal{E}_{2}$ $=k^{2}(q_{1}^{(1)}-q_{2}^{(1)})\int_{\mathcal{S}_{r_{0}}}x_{1}\tilde{\delta}u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}+k^{2}(q_{1}^{(2)}-q_{2}^{(2)})\int_{\mathcal{S}_{r_{0}}}x_{2}\tilde{\delta}u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}.$
@@
$|\mathcal{E}_{2}|$ $\lesssim\int_{\mathcal{S}_{r_{0}}}|\mathbf{x}||\tilde{\delta}u_{2}^{-}||u_{0}|\mathrm{d}\mathbf{x}\lesssim\tau^{-(3+\alpha)}+\tau^{-(\frac{53}{12}+\alpha)}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau}).$ (3.41)

### v3

@@ 3. Unique recovery results for the conductive polygonal-nest and polygonal-cell medium scatterers
$\mathcal{P}_{4}$ $=k^{2}(q_{1}^{(1)}-q_{2}^{(1)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{1}\delta u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}+k^{2}(q_{1}^{(2)}-q_{2}^{(2)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{2}\delta u_{2}^{-}u_{0}\mathrm{d}\mathbf{x},$
@@
$\mathcal{E}_{2}$ $=k^{2}(q_{1}^{(1)}-q_{2}^{(1)})\int_{\mathcal{S}_{r_{0}}}x_{1}\delta u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}+k^{2}(q_{1}^{(2)}-q_{2}^{(2)})\int_{\mathcal{S}_{r_{0}}}x_{2}\delta u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}.$
@@
$|\mathcal{E}_{2}|$ $\lesssim\int_{\mathcal{S}_{r_{0}}}|\mathbf{x}||\delta u_{2}^{-}||u_{0}|\mathrm{d}\mathbf{x}\lesssim\tau^{-(2+\alpha)}+\tau^{-(\frac{41}{12}+\alpha)}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau}).$ (3.40)